A gas turbine operating on a simple Joule cycle has a pressure compression ratio of 8:1. The turbine has a thermal efficiency of 60% of the ideal. Determine the actual thermal efficiency.


1
Expert's answer
2021-05-24T18:46:02-0400

"M_{therm(ideal)}=1-\frac{1}{\frac{g-1}{r^g}}"

"=1-\frac{1}{\frac{1.4-1}{8^{1.4}}}"

"=0.4479"

Actual efficiency of the gas turbine.

"M_{therm(actual)}=0.6\times\eta_{therm(ideal)}"

"=0.6 \times 0.4479"

"=0.2687"

"\eta_{therm(actual)}=26.87%" %